OpenStudy (sweet.teddybear16):

Javier has a job installing windows. He earns a rate of $75 per day plus $8 per window installed. He also receives a stipend of $25 for snacks. Javier knows his weekly earnings can be shown with the following expression: 75d + 8w + 25 Part A: Identify a coefficient, a variable, and a constant in this expression. Part B: If Javier works for 5 days and installs 48 windows, how much does he earn? Show your work to receive full credit. Part C: If Javier gets an increase in his stipend and receives $40 for snacks, would the coefficient, variable, or constant in the equation change? why?

OpenStudy (cas_fangirl_14):

\(\color{Purple}{Part~A}\): coefficient: 75 Variable: d constant: 25 \(\color{Purple}{Part~B}\): 75d+8w+25= 75*5+8*48+25= 375+384+25= 759+25= 784 \(\color{Purple}{Part~C}\): the constant would change, because there is no variable attached to the number.